Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Free Software

Free as in Freedom Linux for Teachers

Nick Urbanik <nicku@nicku.org>

This document Licensed under GPL—see section 5

5, 6 December 2005

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Outline

Why Free Software?

Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Why Free Software?

▶ Okay, I go along with:

▶ saving the whales ▶ feeding the hungry ▶ combatting the ignorance about the evil treatment of

aboriginal people ▶ caring about slaughter of innocent people in Sudan and Iraq ▶ But Free Software? ▶ Is it worth getting all righteous about free software?

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Why Free Software?

▶ Okay, I go along with:

▶ saving the whales ▶ feeding the hungry ▶ combatting the ignorance about the evil treatment of

aboriginal people ▶ caring about slaughter of innocent people in Sudan and Iraq ▶ But Free Software? ▶ Is it worth getting all righteous about free software?

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Why Free Software?

▶ Okay, I go along with:

▶ saving the whales ▶ feeding the hungry ▶ combatting the ignorance about the evil treatment of

aboriginal people ▶ caring about slaughter of innocent people in Sudan and Iraq ▶ But Free Software? ▶ Is it worth getting all righteous about free software?

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Why Free Software?

▶ Okay, I go along with:

▶ saving the whales ▶ feeding the hungry ▶ combatting the ignorance about the evil treatment of

aboriginal people ▶ caring about slaughter of innocent people in Sudan and Iraq ▶ But Free Software? ▶ Is it worth getting all righteous about free software?

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Why Free Software?

▶ Okay, I go along with:

▶ saving the whales ▶ feeding the hungry ▶ combatting the ignorance about the evil treatment of

aboriginal people ▶ caring about slaughter of innocent people in Sudan and Iraq ▶ But Free Software? ▶ Is it worth getting all righteous about free software?

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Why Free Software?

▶ Okay, I go along with:

▶ saving the whales ▶ feeding the hungry ▶ combatting the ignorance about the evil treatment of

aboriginal people ▶ caring about slaughter of innocent people in Sudan and Iraq ▶ But Free Software? ▶ Is it worth getting all righteous about free software?

Free Software Nick Urbanik Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Yes it is!

▶ Bless your computer, my

child! ▶ To join the Church of

Emacs, ▶ you need only say the

Confession of the Faith three times:

▶ There is no system but GNU, and Linux is one of its kernels.

▶ And Bless you too, Saint IGNUtius.

Free Software Nick Urbanik Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Yes it is!

▶ Bless your computer, my

child! ▶ To join the Church of

Emacs, ▶ you need only say the

Confession of the Faith three times:

▶ There is no system but GNU, and Linux is one of its kernels.

▶ And Bless you too, Saint IGNUtius.

Free Software Nick Urbanik Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Yes it is!

▶ Bless your computer, my

child! ▶ To join the Church of

Emacs, ▶ you need only say the

Confession of the Faith three times:

▶ There is no system but GNU, and Linux is one of its kernels.

▶ And Bless you too, Saint IGNUtius.

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Richard Stallman’s Tin Foil Hat

▶ 18 November 2005: Richard Stallman and Bruce Perens were invited by UN as keynote speakers at the UN World Summit on the information society in Tunis, Tunisia

▶ Richard is opposed to RF ID, because of the many privacy

violations possible

▶ A promise was given at the 2003 WSIS in Geneva to not

issue RFIDs again ▶ He was given one on entry

Free Software Nick Urbanik Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Richard Stallman causes Trouble

▶ He wrapped huge amounts of aluminium foil around it,

wearing it very prominently ▶ Richard offered foil to anyone who would take some ▶ He willingly unwrapped it to go through any of the visible

check-points, he simply objected to the potential that people might be reading the RF ID without his knowledge and tracking him around the grounds.

▶ UN security would not let him leave for a long time ▶ Then would not let him in for his next talk, delaying the

whole proceeding by a long time.

Free Software Nick Urbanik Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Sony’s Rootkit

▶ A rootkit is a set of software tools frequently used by

crackers after gaining access to a computer system. ▶ These tools are intended to conceal running processes,

files or system data ▶ which helps an intruder maintain access to a system

without the user’s knowledge. ▶ Sony Corporation have been distributing rootkits on their

music CDs over the last year and a half ▶ The software is installed on your computer without your

permission

▶ even if you decline the EULA

▶ Why? To protect the music industry from piracy. ▶ Sony have ripped off lots of free software to make the root

kit without regard for the license terms, including code written by DVD Jon.

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Poor Celine Lovers!

▶ Foxtrot, 21 November 2005

Free Software Nick Urbanik Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

DVD Jon

▶ The autumn of 1999 Jon Lech Johansen (“DVD Jon”) with others decrypted the copy control system on DVD records to make it possible to view the DVDs at the Linux operating system.

▶ The program that made this possible was called “DeCSS”

and released on the Internet. ▶ This resulted in him being sued by DVD Copy Control

Association in USA, ▶ arrested by Norwegian police and ▶ charged by The Norwegian National Authority for

Investigation and Prosecution of Economic and Environmental Crime (ØKOKRIM).

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

▶

Free Software Nick Urbanik Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Resources

Sam Williams. Free as in Freedom: Richard Stallman’s Crusade for Free Software. O’Reilly 2002.

http://oreilly.com/openbook/freedom/

Bruce Perens. Friday, November 18: Richard Stallman Gets in Trouble with UN Security for Wearing a Tin-Foil Hat.

http://perens.sourcelabs.com/
http://lwn.net/Articles/160913/

Jon Lech Johansen. So Sue Me!

http://nanocrew.net/
http://efn.no/jonjohansen-en.html

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Topics Covered

Why Free Software?

Richard Stallman’s Tin Foil Hat

Sony’s Rootkit

DVD Jon

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Topics Covered

Why Free Software?

Richard Stallman’s Tin Foil Hat

Sony’s Rootkit

DVD Jon

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Topics Covered

Why Free Software?

Richard Stallman’s Tin Foil Hat

Sony’s Rootkit

DVD Jon

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Topics Covered

Why Free Software?

Richard Stallman’s Tin Foil Hat

Sony’s Rootkit

DVD Jon

Free Software Nick Urbanik

Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

Topics Covered

Why Free Software?

Richard Stallman’s Tin Foil Hat

Sony’s Rootkit

DVD Jon

Free Software Nick Urbanik Why Free Software? Richard Stallman’s Tin Foil Hat Sony’s Rootkit DVD Jon License Of This Document

License Of This Document

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005 Nick Urbanik <nicku@nicku.org> You can redistribute modified or unmodified copies of this document provided that this copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation—either version 2 of the License or (at your option) any later version.